# An Axiomatic Risk-Reward Framework for Sustainable Investing

### v3

@@ 3.3.1 ESG-AVaR
We demonstrate the two approaches presented in equations (13) and (14) to develop ESG-coherent risk measures based on AVaR. Given a random bivariate vector $X_{T}=[r_{T},\;{ESG}_{T}]^{\prime}\in\mathcal{X}_{2}=\mathcal{L}_{1}(\Omega,\mathcal{F},P;\mathbb{R}^{2})$, the first measure, given by (13), is171717 See Ogryczak and Ruszczynski, (2002); Rockafellar and Uryasev, (2002) for the extremal representation.
${ESG}\text{-}\text{AVaR}_{\lambda,\tau}\left(\left[\begin{matrix}r_{T}\\ {ESG}_{T}\end{matrix}\right]\right)$ $:=\text{AVaR}_{\tau}\left((1-\lambda)r_{T}+\lambda{ESG}_{T}\right)$
@@
where $(a)^{+}$ denotes max(a,0)(a,0). Following the discussion on (13), we conclude that ${ESG}\text{-}\text{AVaR}_{\lambda,\tau}$ is coherent. It is similar to the multivariate expected shortfall presented by Ekeland et al., (2012) (although their measure lacks the parametrization using $\lambda$). Since ${ESG}\text{-}\text{AVaR}_{\lambda,\tau}(\cdot)$ is computed on univariate data (i.e., as a linear combination of $r$ and ESG{ESG}), numerical applications using ${ESG}\text{-}\text{AVaR}_{\lambda,\tau}(\cdot)$ do not present any particular challenge; it is possible to fully utilize existing procedures developed for AVaR for risk estimation, portfolio optimization, and risk management. (See, e.g. Shapiro et al.,, 2021.)
The dual representation of ${ESG}\text{-}\text{AVaR}_{\lambda,\tau}(X_{T})$ is
${ESG}\text{-}\text{AVaR}_{\lambda,\tau}\left(\left[\begin{matrix}r_{T}\\ {ESG}_{T}\end{matrix}\right]\right)=\sup_{[\zeta_{1},\;\zeta_{2}]^{\prime}\in\mathcal{A}_{{ESG}\text{-}\text{AVaR}_{\lambda,\tau}(X_{T})}}\left(-\mathbb{E}[r_{T}\zeta_{1}+{ESG}_{T}\zeta_{2}]\right),$ (16)
@@
$\mathcal{A}_{{ESG}\text{-}\text{AVaR}_{\lambda,\tau}}=\left\{[\zeta_{1},\;\zeta_{2}]^{\prime}\in\mathcal{L}_{\infty}(\Omega,\mathcal{F},P;\mathbb{R}^{2}):\;[\zeta_{1},\;\zeta_{2}]^{\prime}=\xi[1-\lambda,\;\lambda]^{\prime};\;0\leq\xi\leq\frac{1}{1-\tau}\;\text{a.s. };\mathbb{E}[\xi]=1\right\}.$ (17)
@@
${ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}\left(\left[\begin{matrix}r_{T}\\ {ESG}_{T}\end{matrix}\right]\right)$ $:=(1-\lambda)\text{AVaR}_{\tau}(r_{T})+\lambda\text{AVaR}_{\tau}({ESG}_{T}).$ (18)
${ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}(\cdot)$ is also ESG-coherent. It can be viewed as the limit of ${ESG}\text{-}\text{AVaR}_{\lambda,\tau}(\cdot)$ in the case of an asset for which it is not possible to diversify between the monetary and ESG components as they are comonotone. From an economic perspective, ${ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}(\cdot)$ is significant for investors who consider the worst-case scenario in terms of the dependency structure.
The dual representation of ${ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}(X_{T})$ is
${ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}\left(\left[\begin{matrix}r_{T}\\ {ESG}_{T}\end{matrix}\right]\right)=\sup_{[\zeta_{1},\;\zeta_{2}]^{\prime}\in\mathcal{A}_{{ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}(X_{T})}}-\mathbb{E}[r_{T}\zeta_{1}+{ESG}_{T}\zeta_{2}],$ (19)
@@
$\mathcal{A}_{{ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}}=\left\{[\zeta_{1},\;\zeta_{2}]^{\prime}\in\mathcal{L}_{\infty}(\Omega,\mathcal{F},P;\mathbb{R}^{2}):\mathbb{E}[\zeta_{1}]=1-\lambda;\mathbb{E}[\zeta_{2}]=\lambda;\zeta_{1},\zeta_{2}\geq 0;\zeta_{1}\leq\frac{1-\lambda}{1-\tau};\zeta_{2}\leq\frac{\lambda}{1-\tau}\right\}.$ (20)
The derivation of the dual representation of ${ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}(\cdot)$ is also given in Appendix C.
@@ 3.3.2 Non-ESG-coherent measure examples
It is well known that the standard deviation $\sigma$ (the volatility) and the variance $\mathbb{V}$ are not coherent risk measures. We consider the application of (13) and (14) to $\sigma$ and $\mathbb{V}$ and show that, in all cases, the result is an ESG measure that does not satisfy the ESG-coherency axioms.
@@
Table 1: Summary of the properties satisfied by ESG risk measures.
@@
${ESG}\text{-}\text{AVaR}_{\lambda,\tau}$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
${ESG}\text{-}\text{AVaR}_{\lambda,\tau}^{l}$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ 4 ESG reward–risk ratios
It is natural to extend the ESG framework to reward–risk ratios (RRRs), used to measure risk-adjusted performance of an investment. Following Cheridito and Kromer, (2013), a reward–risk ratio $\alpha(r)$ in a univariate setting is
@@
where $\theta(r_{T}):\mathcal{X}\rightarrow\mathbb{R}\cup\{\pm\infty\}$ and $\rho(r_{T}):\mathcal{X}\rightarrow\mathbb{R}\cup\{+\infty\}$ are reward and risk measures, respectively. Cheridito and Kromer, (2013) identified four conditions desirable for RRRs:
@@
where $\boldsymbol{\theta}_{\lambda}(X_{T})$ and $\boldsymbol{\rho}_{\lambda}(X_{T})$ are ESG reward and risk measures as defined in Section 3 and Appendix A.181818 In principle, it is possible for the risk and reward measures to have different values of $\lambda$. We consider the case of values of lambda for both the numerator and the denominator for conciseness. The extension of the Cheridito-Kromer conditions for ESG-RRRs are:
@@
Verification of DB-RM depends on the bivariate distribution of $X_{T}\in\mathcal{X}_{2}$ and not on the univariate distribution of the returns. As in the case of the risk measure axiom MO-M, verification of MO-RM requires the use of partial ordering.
@@
In general, risk and reward measures may not depend on the distribution of returns and ESG under a single probability measure, as in the case of robust reward–risk ratios, which take into account the fact that agents do not know with certainty the distribution of the random variables (Cheridito and Kromer,, 2013).
@@ 4.1 Examples of ESG reward–risk ratios
We present six examples of ESG reward–risk ratios derived from RRRs commonly used in the literature. The ratios are obtained by generalizing the univariate reward–risk ratios using the approach described by (13). Note that, as in the case of risk measures, the definition of a reward–risk ratio can be based on several alternative specifications of the random variable $X_{T}=[r_{T},\;{ESG}_{T}]^{\prime}$. In particular, the ratios can be computed using the returns, the excess returns over a risk-free rate, the final wealth, profit and/or losses, etc. The same logic applies to the ESG component. Here, we only provide hints concerning which approach to use in practice; the choice depends on the specific needs of the practitioner or regulator who uses these measures.
@@
    $\text{ESG-}\text{SR}_{\lambda}(X_{T}):=\dfrac{\text{ESG-}\mu_{\lambda}(X_{T}-\text{SA}_{T})}{\text{ESG-}\sigma_{\lambda}(X_{T})}.$ (27)
@@
    the ESG standard deviation $\text{ESG-}\sigma_{\lambda}(X_{T})$ is given by (22), and $\text{SA}_{T}=[\text{RF}_{T}^{r},\;\text{RF}_{T}^{{ESG}}]^{\prime}$ is the bivariate return of an ESG safe asset. $\text{ESG-}\mu_{\lambda}$ is an ESG-coherent reward measure, while $\text{ESG-}\sigma_{\lambda}$ is not an ESG-coherent risk measure. The ESG Sharpe ratio satisfies QC-RM, SI-RM, and DB-RM, but it does not satisfy condition MO-RM.191919 The SI-RM property applies if $\text{SA}=[0,\;0]^{\prime}$ or if $X$ is intended to represent a vector of excess returns over the risk-free rate.
  • ESG Rachev ratio. The univariate Rachev ratio is defined as (Biglova et al.,, 2004)
@@
    We generalize to an ESG-RR by replacing AVaR with ESG-AVaR. Let $X_{T}\in\mathcal{X}_{2}$. Then,
@@
    Note that ESG-RR satisfies MO-RM, SI-RM and DB-RM but fails to satisfy QC-RM since the numerator is not concave.
@@
  • ESG Omega ratio. Defining $Y_{T}=(1-\lambda)r_{T}+\lambda{ESG}_{T}$, and $F(y)$ as the cumulative distribution function of $Y_{T}$, the ESG version of the Omega ratio (Keating and Shadwick,, 2002) is
@@
    Equivalently (see Farinelli and Tibiletti,, 2008),
@@
    where $m,n\in\mathbb{R}$ and $p,q>0$. This ratio satisfies MO-RM and DB-RM. It also satisfies SI-RM if $n=m=0$. It does not satisfy QC-RM (see the example in Cheridito and Kromer,, 2013, Section 4.3).
Table 2 summarizes the properties of these six ESG-RRRs.
Table 2: Summary of the properties of ESG reward–risk ratios.
@@ 5 Empirical analysis
We present an empirical analysis in which we estimate a set of daily ESG risk measures and ESG reward–risk ratios. We use such measures to rank equity assets from the Dow-Jones Industrial Average (DJIA) index, assessing the role of $\lambda$ for different measures. We use the log-returns computed on adjusted close prices downloaded from Factset, and for the ESG component we consider the Factset Truvalue scores. These ESG scores, in contrast to scores by other data providers, are not computed on the basis of balance sheet items and metrics reported by companies, but instead they are obtained by analysing news and documents, performing sentiment analyses on news and event tracking on sustainability-related events with artificial intelligence techniques. Such scores are based on a large number of sources, and reflect the public perception of the investors on the sustainability of companies. One unique feature of these scores is that they are updated with daily frequency, allowing the dynamics of the sustainability of each company to be tracked. In contrast, other data providers update ESG scores with quarterly or yearly frequency. The Truvalue scores represent therefore a desirable dataset for showcasing ESG-coherent risk measures and reward ratios. Truvalue scores are provided in two variants: Pulse Scores and Insight Scores, where the former focus more on near-term performance changes, and the latter considers the longer-term track record. In this analysis we consider Pulse Scores, as they provide a more dynamic measure of the sustainability profile of the companies.
@@
We consider 28 constituents of the DJIA index as of December 31, 2023. We use the daily time series of log-returns ($r_{T}$) and Truvalue Pulse Scores (${ESG}_{T}$) in the period from January 1, 2020 to December 31, 2023. The daily Pulse Score is normalized between 1/c-1/c and $1/c$, where $c=252$ (the number of trading days in a year). [^21] The normalization leads to a broadly similar scale of average returns and average ESG (see Table 3). The scaling of the ESG score, as long as it is consistent across assets, does not affect the generality of the results, as it can be compensated by changing the scaling of the parameter $\lambda$.
@@
Name Ticker GICS Sector Name $\mathbb{E}[r]$ $\sigma(r)$ $\mathbb{E}[{ESG}]$ $\sigma({ESG})$ $\rho(r,{ESG})$
Apple Inc. AAPL Information Technology $0.242$ $0.336$ 0 $0.005$ 0.025-0.025
Amgen Inc. AMGN Health Care $0.045$ $0.262$ $0.149$ $0.012$ 0.002-0.002
American Express Co. AXP Financials $0.102$ $0.409$ $0.15$ $0.01$ 0.015-0.015
Boeing Company BA Industrials 0.056-0.056 $0.546$ 0.002-0.002 $0.009$ $0.041$
Caterpillar Inc. CAT Industrials $0.174$ $0.34$ $0.376$ $0.006$ 0.019-0.019
Salesforce, Inc. CRM Information Technology $0.121$ $0.407$ $0.275$ $0.007$ 0.035-0.035
Cisco Systems, Inc. CSCO Information Technology $0.013$ $0.293$ $0.18$ $0.008$ $0.06$
Chevron Corporation CVX Energy $0.053$ $0.39$ $0.117$ $0.007$ $0.018$
Walt Disney Company DIS Communication Services 0.118-0.118 $0.357$ 0.03-0.03 $0.008$ $0.035$
Goldman Sachs Gr., Inc. GS Financials $0.13$ $0.347$ 0.056-0.056 $0.01$ 0.03-0.03
Home Depot, Inc. HD Consumer Discretionary $0.116$ $0.31$ $0.029$ $0.01$ 0.008-0.008
Honeywell Int. Inc. HON Industrials $0.042$ $0.29$ $0.505$ $0.007$ $0.047$
IBM Corporation IBM Information Technology $0.061$ $0.278$ $0.246$ $0.005$ $0.011$
Intel Corporation INTC Information Technology 0.044-0.044 $0.415$ $0.253$ $0.008$ $0.007$
Johnson & Johnson JNJ Health Care $0.018$ $0.206$ 0.179-0.179 $0.008$ $0.024$
JPMorgan Chase & Co. JPM Financials $0.05$ $0.344$ 0.047-0.047 $0.009$ 0.022-0.022
Coca-Cola Company KO Consumer Staples $0.016$ $0.225$ $0.322$ $0.004$ $0.037$
McDonald’s Corp. MCD Consumer Discretionary $0.102$ $0.247$ $0.125$ $0.006$ 0.01-0.01
3M Company MMM Industrials 0.12-0.12 $0.286$ 0.02-0.02 $0.013$ $0.01$
Merck & Co., Inc. MRK Health Care $0.057$ $0.241$ $0.118$ $0.005$ $0.023$
Microsoft Corporation MSFT Information Technology $0.218$ $0.326$ $0.015$ $0.006$ $0.011$
NIKE, Inc. Class B NKE Consumer Discretionary $0.017$ $0.354$ $0.1$ $0.008$ 0.007-0.007
Procter & Gamble Co. PG Consumer Staples $0.04$ $0.222$ $0.303$ $0.009$ 0.024-0.024
Travelers Companies, Inc. TRV Financials $0.083$ $0.313$ 0.124-0.124 $0.014$ 0.021-0.021
UnitedHealth Group Inc. UNH Health Care $0.146$ $0.306$ $0.042$ $0.008$ $0.016$
Verizon Comm. Inc. VZ Communication Services 0.122-0.122 $0.217$ $0.162$ $0.007$ $0.005$
Walgreens Boots All., Inc. WBA Consumer Staples 0.204-0.204 $0.361$ 0.19-0.19 $0.013$ $0.017$
Walmart Inc. WMT Consumer Staples $0.071$ $0.237$ $0.074$ $0.006$ 0.006-0.006
@@
Table 3 reports, by company, the distribution mean and standard deviation of daily returns and ESG scores, as well as the correlation between returns and ESG scores. The dataset presents a wide cross-sectional variability in terms of the average return and risk, as well as in terms of the average ESG scores. ESG data show a smaller volatility compared to the returns despite a comparable cross-sectional dispersion of average returns and average ESG. This reflects the fact that the changes in the sustainability profile of a company happen at a much slower pace than the shifts seen in the monetary value of a stock. The correlation between the daily returns and ESG score is very close to zero for all the stocks. Figure 2 displays the correlation matrix between the returns and ESG of each stock (the first 28 rows and columns are the returns, the latter 28 are the ESG values). We see that the returns show consistently positive and strong correlation with each other. In contrast, ESG-to-ESG correlations have mixed signs. ESG-to-return correlations are very close to zero. In terms of risk management, the lack of positive correlations among ESG and returns is beneficial to investors, as it allows diversification of risk between the two.

### v4

@@ 3.3.1 ESG-AVaR
We demonstrate the two approaches presented in equations (13) and (14) to develop ESG-coherent risk measures based on AVaR. Given a random bivariate vector $X_{T}=[r_{T},\;{ESG}_{T}]^{\prime}\in\mathcal{X}_{2}=\mathcal{L}_{1}(\Omega,\mathcal{F},P;\mathbb{R}^{2})$, the first measure, given by (13), is [^16] See Ogryczak and Ruszczyński (2002); Rockafellar and Uryasev (2002) for the extremal representation.
$\text{ESG-}\text{AVaR}_{\lambda,\tau}\left(\left[\begin{matrix}r_{T}\\ {ESG}_{T}\end{matrix}\right]\right)$ $:=\text{AVaR}_{\tau}\left((1-\lambda)r_{T}+\lambda{ESG}_{T}\right)$
@@
where $(a)^{+}$ denotes max(a,0)(a,0). Following the discussion on (13), we conclude that $\text{ESG-}\text{AVaR}_{\lambda,\tau}$ is coherent. It is similar to the multivariate expected shortfall presented by Ekeland et al. (2012) (although their measure lacks the parametrization using $\lambda$ allowing investor-specific ESG preferences). For $\lambda=0$ the measure reduces to the univariate AVaR. Since $\text{ESG-}\text{AVaR}_{\lambda,\tau}(\cdot)$ is computed on univariate data (i.e., as a linear combination of $r$ and ESG{ESG}), numerical applications using $\text{ESG-}\text{AVaR}_{\lambda,\tau}(\cdot)$ do not present any particular challenge; it is possible to fully utilize existing procedures developed for AVaR for risk estimation, portfolio optimization, and risk management. (See, e.g. Shapiro et al., 2021).
The dual representation of $\text{ESG-}\text{AVaR}_{\lambda,\tau}(X_{T})$ is
$\text{ESG-}\text{AVaR}_{\lambda,\tau}\left(\left[\begin{matrix}r_{T}\\ {ESG}_{T}\end{matrix}\right]\right)=\sup_{[\zeta_{1},\;\zeta_{2}]^{\prime}\in\mathcal{A}_{\text{ESG-}\text{AVaR}_{\lambda,\tau}(X_{T})}}\left(-\mathbb{E}[r_{T}\zeta_{1}+{ESG}_{T}\zeta_{2}]\right),$ (16)
@@
$\mathcal{A}_{\text{ESG-}\text{AVaR}_{\lambda,\tau}}=\left\{[\zeta_{1},\;\zeta_{2}]^{\prime}\in\mathcal{L}_{\infty}(\Omega,\mathcal{F},P;\mathbb{R}^{2}):\;[\zeta_{1},\;\zeta_{2}]^{\prime}=\xi[1-\lambda,\;\lambda]^{\prime};\;0\leq\xi\leq\frac{1}{1-\tau}\;\text{a.s. };\mathbb{E}[\xi]=1\right\}.$ (17)
@@
$\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}\left(\left[\begin{matrix}r_{T}\\ {ESG}_{T}\end{matrix}\right]\right)$ $:=(1-\lambda)\text{AVaR}_{\tau}(r_{T})+\lambda\text{AVaR}_{\tau}({ESG}_{T}).$ (18)
$\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}(\cdot)$ is also ESG-coherent. It can be viewed as the limit of $\text{ESG-}\text{AVaR}_{\lambda,\tau}(\cdot)$ in the case of an asset for which it is not possible to diversify between the monetary and ESG components as they are comonotone. From an economic perspective, $\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}(\cdot)$ is significant for investors who consider the worst-case scenario in terms of the dependency structure.
The dual representation of $\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}(X_{T})$ is
$\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}\left(\left[\begin{matrix}r_{T}\\ {ESG}_{T}\end{matrix}\right]\right)=\sup_{[\zeta_{1},\;\zeta_{2}]^{\prime}\in\mathcal{A}_{\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}(X_{T})}}-\mathbb{E}[r_{T}\zeta_{1}+{ESG}_{T}\zeta_{2}],$ (19)
@@
$\mathcal{A}_{\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}}=\left\{[\zeta_{1},\;\zeta_{2}]^{\prime}\in\mathcal{L}_{\infty}(\Omega,\mathcal{F},P;\mathbb{R}^{2}):\mathbb{E}[\zeta_{1}]=1-\lambda;\mathbb{E}[\zeta_{2}]=\lambda;\zeta_{1},\zeta_{2}\geq 0;\zeta_{1}\leq\frac{1-\lambda}{1-\tau};\zeta_{2}\leq\frac{\lambda}{1-\tau}\right\}.$ (20)
The derivation of the dual representation of $\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}(\cdot)$ is also given in Appendix C.
@@ 3.3.2 Non-ESG-coherent measure examples
In order to better delineate the boundaries of the axiomatic framework, we discuss here examples of measures that are not ESG-coherent. It is well known that the standard deviation $\sigma$ (the volatility) and the variance $\mathbb{V}$ are not coherent risk measures. We consider the application of (13) and (14) to $\sigma$ and $\mathbb{V}$ and show that, in all cases, the result is an ESG measure that does not satisfy the ESG-coherency axioms.
@@
Table 1: Summary of ESG risk measures and their properties. A checkmark indicates that the corresponding property is satisfied.
@@
$\text{ESG-}\text{AVaR}_{\lambda,\tau}$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
$\text{ESG-}\text{AVaR}_{\lambda,\tau}^{l}$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ 4 ESG reward–risk ratios
It is natural to extend the ESG framework to reward–risk ratios (RRRs), used to measure risk-adjusted performance of an investment. Following Cheridito and Kromer (2013), a reward–risk ratio $\alpha(r)$ in a univariate setting is
@@
where $\theta(r_{T}):\mathcal{X}\rightarrow\mathbb{R}\cup\{\pm\infty\}$ and $\rho(r_{T}):\mathcal{X}\rightarrow\mathbb{R}\cup\{+\infty\}$ are reward and risk measures, respectively. Cheridito and Kromer (2013) identified four conditions desirable for RRRs:
@@
where $\boldsymbol{\theta}_{\lambda}(X_{T})$ and $\boldsymbol{\rho}_{\lambda}(X_{T})$ are ESG reward and risk measures as defined in Section 3 and Appendix A using the same preference parameter $\lambda$. [^17] In principle, it is possible for the risk and reward measures to have different values of $\lambda$. We consider the case of values of lambda for both the numerator and the denominator for conciseness. The extension of the Cheridito-Kromer conditions for ESG-RRRs are:
@@
Verification of DB-RM depends on the bivariate distribution of $X_{T}\in\mathcal{X}_{2}$ and not on the univariate distribution of the returns.
@@
Proposition 5 does not guarantee that the DB-RM property is satisfied since, as discussed in (Cheridito and Kromer, 2013), risk and reward measures may not depend on the distribution of returns and ESG under a single probability measure, and in such case the distribution-based property is not satisfied. This is for example the case of distributionally robust reward–risk ratios, which consider the fact that agents do not know with certainty the distribution of the random variables, and consider a whole family of distributions (see e.g. Liu et al., 2017).
@@ 4.1 Examples of ESG reward–risk ratios
We present six examples of ESG reward–risk ratios derived from RRRs commonly used in the literature. The ratios are obtained by generalizing the univariate reward–risk ratios using the approach described by (13). Note that, as in the case of risk measures, the definition of a reward–risk ratio can be based on several alternative specifications of the random variable $X_{T}=[r_{T},\;{ESG}_{T}]^{\prime}$. In particular, the ratios can be computed using the returns, the excess returns over a risk-free rate, the final wealth, profit and/or losses, etc. The same logic applies to the ESG component. Here, we only provide illustrative guidance concerning which approach to use in practice; the choice depends on the specific needs of the practitioner or regulator who uses these measures.
@@
    $\text{ESG-}\text{SR}_{\lambda}(X_{T}):=\dfrac{\text{ESG-}\mu_{\lambda}(X_{T}-\text{SA}_{T})}{\text{ESG-}\sigma_{\lambda}(X_{T})},$ (27)
@@
    the ESG standard deviation $\text{ESG-}\sigma_{\lambda}(X_{T})$ is given by (22), and $\text{SA}_{T}=[\text{RF}_{T}^{r},\;\text{RF}_{T}^{{ESG}}]^{\prime}$ is the bivariate return of an ESG safe asset. $\text{ESG-}\mu_{\lambda}$ is an ESG-coherent reward measure, while $\text{ESG-}\sigma_{\lambda}$ is not an ESG-coherent risk measure. The ESG Sharpe ratio satisfies QC-RM, SI-RM, and DB-RM, but it does not satisfy condition MO-RM. [^18] The SI-RM property applies if $\text{SA}=[0,\;0]^{\prime}$ or if $X$ is intended to represent a vector of excess returns over the risk-free rate.
  • ESG Rachev ratio. The univariate Rachev ratio is defined as (Biglova et al., 2004)
@@
    We generalize to an ESG-RR by replacing AVaR with ESG-AVaR. Let $X_{T}\in\mathcal{X}_{2}$. Then
@@
    Note that ESG-RR satisfies MO-RM, SI-RM and DB-RM but, as in the univariate case, fails to satisfy QC-RM since the numerator is not concave.
@@
  • ESG Omega ratio. Defining $Y_{T}=(1-\lambda)r_{T}+\lambda{ESG}_{T}$, and $F(y)$ as the cumulative distribution function of $Y_{T}$, the ESG version of the Omega ratio (Keating and Shadwick, 2002) is
@@
    Equivalently (see Farinelli and Tibiletti, 2008),
@@
    where $m,n\in\mathbb{R}$ and $p,q>0$. This ratio satisfies MO-RM and DB-RM. It also satisfies SI-RM if $n=m=0$. Mirroring the univariate case, it does not satisfy QC-RM (see the example in Cheridito and Kromer, 2013, Section 4.3).
Table 2 summarizes the properties satisfied by each of these six ratios.
Table 2: Summary of ESG reward–risk ratios and their properties. A checkmark indicates that the corresponding property is satisfied.
@@ 5 Empirical analysis
We present an empirical analysis in which we estimate a set of daily ESG risk measures and ESG reward–risk ratios. We use such measures to rank equity assets from the Dow-Jones Industrial Average (DJIA) index, and the main goal is to illustrate how rankings vary with the preference parameter $\lambda$. We use the log-returns computed on adjusted close prices downloaded from Factset, and for the ESG component we consider the Factset Truvalue scores. These ESG scores, in contrast to scores by other data providers, are not computed on the basis of balance sheet items and metrics reported by companies, but instead they are obtained by analysing news and documents, performing sentiment analyses on news and event tracking on sustainability-related events with artificial intelligence techniques. Such scores are based on a large number of sources, and reflect the public perception of the investors on the sustainability of companies. One unique feature of these scores is that they are updated with daily frequency, allowing the dynamics of the sustainability of each company to be tracked. In contrast, other data providers update ESG scores with quarterly or yearly frequency. The Truvalue scores represent therefore a desirable dataset for showcasing ESG-coherent risk measures and reward ratios. Truvalue scores are provided in two variants: Pulse Scores and Insight Scores, where the former focus more on near-term performance changes, and the latter considers the longer-term track record. In this analysis we consider Pulse Scores, as they provide a more dynamic measure of the sustainability profile of the companies. The high-frequency of the Truvalue dataset however comes at the cost of a less transparent scoring methodology compared to competitors due to the reliance to AI techniques.
@@
We consider 28 constituents of the DJIA index as of December 31, 2023. We use the daily time series of log-returns ($r_{T}$) and Truvalue Pulse Scores (${ESG}_{T}$) in the period from January 1, 2020 to December 31, 2023. The daily Pulse Score is normalized between 1/c-1/c and $1/c$, where $c=252$ (the number of trading days in a year). The normalization leads to a broadly similar scale of average returns and average ESG (see Table 3) The scaling of the ESG score, as long as it is consistent across assets, does not affect the generality of the results, as it can be compensated by changing the scaling of the subjective parameter $\lambda$. One concern with the dataset is that, when ESG scores are strictly bounded, some axiomatic conditions may lose their standard economic interpretation. In applications where this issue is relevant, one may adopt the modeling convention of working with a monotone transformation of the scores that has unbounded support, thereby restoring a more formal setting without altering their ordinal content.
@@
Name GICS Sector $\mathbb{E}[r]$ $\sigma(r)$ $\mathbb{E}[{ESG}]$ $\sigma({ESG})$ $\rho(r,{ESG})$
Apple Inc. IT $0.242$ $0.336$ $0.000$ $0.005$ 0.025-0.025
Amgen Inc. Health Care $0.045$ $0.262$ $0.149$ $0.012$ 0.002-0.002
American Express Co. Financials $0.102$ $0.409$ $0.150$ $0.010$ 0.015-0.015
Boeing Company Industrials 0.056-0.056 $0.546$ 0.002-0.002 $0.009$ $0.041$
Caterpillar Inc. Industrials $0.174$ $0.340$ $0.376$ $0.006$ 0.019-0.019
Salesforce, Inc. IT $0.121$ $0.407$ $0.275$ $0.007$ 0.035-0.035
Cisco Systems, Inc. IT $0.013$ $0.293$ $0.180$ $0.008$ $0.060$
Chevron Corporation Energy $0.053$ $0.390$ $0.117$ $0.007$ $0.018$
Walt Disney Company Comm. Services 0.118-0.118 $0.357$ 0.030-0.030 $0.008$ $0.035$
Goldman Sachs Gr., Inc. Financials $0.130$ $0.347$ 0.056-0.056 $0.010$ 0.030-0.030
Home Depot, Inc. Consumer Discr. $0.116$ $0.310$ $0.029$ $0.010$ 0.008-0.008
Honeywell Int. Inc. Industrials $0.042$ $0.290$ $0.505$ $0.007$ $0.047$
IBM Corporation IT $0.061$ $0.278$ $0.246$ $0.005$ $0.011$
Intel Corporation IT 0.044-0.044 $0.415$ $0.253$ $0.008$ $0.007$
Johnson & Johnson Health Care $0.018$ $0.206$ 0.179-0.179 $0.008$ $0.024$
JPMorgan Chase & Co. Financials $0.050$ $0.344$ 0.047-0.047 $0.009$ 0.022-0.022
Coca-Cola Company Consumer Staples $0.016$ $0.225$ $0.322$ $0.004$ $0.037$
McDonald’s Corp. Consumer Discr. $0.102$ $0.247$ $0.125$ $0.006$ 0.010-0.010
3M Company Industrials 0.120-0.120 $0.286$ 0.020-0.020 $0.013$ $0.010$
Merck & Co., Inc. Health Care $0.057$ $0.241$ $0.118$ $0.005$ $0.023$
Microsoft Corporation IT $0.218$ $0.326$ $0.015$ $0.006$ $0.011$
NIKE, Inc. Class B Consumer Discr. $0.017$ $0.354$ $0.100$ $0.008$ 0.007-0.007
Procter & Gamble Co. Consumer Staples $0.040$ $0.222$ $0.303$ $0.009$ 0.024-0.024
Travelers Companies, Inc. Financials $0.083$ $0.313$ 0.124-0.124 $0.014$ 0.021-0.021
UnitedHealth Group Inc. Health Care $0.146$ $0.306$ $0.042$ $0.008$ $0.016$
Verizon Comm. Inc. Comm. Services 0.122-0.122 $0.217$ $0.162$ $0.007$ $0.005$
Walgreens Boots All., Inc. Consumer Staples 0.204-0.204 $0.361$ 0.190-0.190 $0.013$ $0.017$
Walmart Inc. Consumer Staples $0.071$ $0.237$ $0.074$ $0.006$ 0.006-0.006
Table 3 reports, by company, the distribution mean and standard deviation of daily returns and ESG scores, as well as the correlation between returns and ESG scores. The dataset presents a wide cross-sectional variability in terms of the average return and risk, as well as in terms of the average ESG scores. ESG data show a smaller volatility compared to the returns despite a comparable cross-sectional dispersion of average returns and average ESG. This reflects the fact that the changes in the sustainability profile of a company happen at a much slower pace than the shifts seen in the monetary value of a stock. The correlation between the daily returns and ESG score is very close to zero for all the stocks. Figure 2 displays the correlation matrix between the returns and ESG of each stock (the first 28 rows and columns are the returns, the latter 28 are the ESG values). We see that the returns show consistently positive and strong correlation with each other. In contrast, ESG-to-ESG correlations have mixed signs. ESG-to-return correlations are very close to zero. In terms of risk management, the lack of positive correlations among ESG and returns is beneficial to investors, as it allows diversification of risk between the two, supporting the motivation of the axiomatic framework.